OpenStudy (anonymous):

SOLVE: 3x - 2y = 9 and -3x + 2y = -9 Using elimination

OpenStudy (dls):

multiply equation 1by -1 and add both

OpenStudy (jhannybean):

Hmm.... That would give you -3x + 2y = -9 -3x + 2y = -9 ----------------- -9x + 4y = -18 Then you can alter the second equation to be 9x - 6y = 27 by multiplying equation (-3x +2y = -9) by (-3) -9x + 4y = -18 9x - 6y = +27 ------------------ 0x - 2y = 9 -2y = 9 y = -9/2 Substitute y into one of two equations to solve for x.

OpenStudy (dls):

lol i meant multiply 2nd equation by -1

OpenStudy (dls):

-4y=18 y=-9/2 directly

OpenStudy (jhannybean):

:P

OpenStudy (jhannybean):

Either method works xD

OpenStudy (dls):

yup